OpenStudy (anonymous):

What two inequalities would you use to solve the absolute value inequality |4x + 3| > 2?

OpenStudy (igreen):

We split it into two inequalities, the one it is now: 4x + 3 > 2 And we multiply -1 to all the terms inside the absolute inequality brackets.

OpenStudy (igreen):

-1 * 4x = ? -1 * 3 = ?

OpenStudy (igreen):

What do you get?

OpenStudy (anonymous):

negative 3

OpenStudy (igreen):

Yes, and what's -1 * 4x?

OpenStudy (anonymous):

-4x

OpenStudy (anonymous):

these are my answer choices A. 4x + 3 > -2, 4x + 3 < 2 B. 4x + 3 < -2, 4x + 3 > 2 C. 4x + 3 > 2, 4x – 3 < -2 D. 4x + 3 > 2, 4x + 3 > -2

OpenStudy (texaschic101):

4x + 3 > 2 -(4x + 3) > 2 = -4x - 3 > 2 = 4x + 3 < -2

OpenStudy (igreen):

Yes, so we have: -4x - 3 > 2 and 4x + 3 > 2

OpenStudy (igreen):

Which..isn't a choice..lol.

OpenStudy (igreen):

I guess you can divide -1 to both sides too

OpenStudy (texaschic101):

found that out....you have to multiply by -1 and change inequality sign

OpenStudy (igreen):

You mean divide..since it's being multiplied :P

OpenStudy (igreen):

But either way we still switch the sign, since it's negative.

OpenStudy (texaschic101):

you can still multiply by -1....either way will work